THE B^{3}Σ_{u}^{−} ← b^{1}Σ_{g}^{+} TRANSITION OF MOLECULAR OXYGEN

B. R. LEWIS, S. T. GIBSON, Research School of Physical Sciences and Engineering, The Australian National University, Canberra, ACT 0200, Australia; T. G. SLANGER, \underline{D. L. HUESTIS}, Molecular Physics Laboratory, SRI International, Menlo Park, CA 94025.

The origin of the intensity of the B^{3}Σ_{u}^{−} ← b^{1}Σ_{g}^{+} transition of molecular oxygen, first observed recently by Eppink et al. [J. Chem. Phys. 108, 1305 (1998).], is discussed. It is shown that the B ← b transition borrows its intensity principally from the dipole-allowed B ← X transition, through spin-orbit mixing between the X^{3}Σ_{g}^{−} and b^{1}Σ_{g}^{+} states. Estimated continuum photoabsorption cross sections and discrete oscillator strengths for the B ← b system are presented.